# SuperARC: Can Increasing Complexity Explain Intelligence? A Test for Artificial Super Intelligence Based On the Principles of Causal Recursive Compression and Algorithmic Probability

### v4

@@ Abstract
We introduce an open-ended test grounded in Kolmogorov-Chaitin complexity, information theory, and algorithmic probability that can avoid benchmark contamination in the quantitative evaluation of frontier models in the context of their Artificial General Intelligence (AGI) and Superintelligence (ASI) claims. Unlike other tests, this test does not rely on statistical compression methods (such as GZIP or LZW), which are more closely related to Shannon entropy than to Kolmogorov-Chaitin complexity and are not able to test beyond simple pattern matching. The test challenges aspects of AI, in particular LLMs, related to features of intelligence of fundamental nature such as synthesis and model creation in the context of inverse problems (generating new knowledge from observation). We argue that metrics based on model abstraction and abduction (optimal Bayesian ‘inference’) for predictive ‘planning’ can provide a robust framework for testing intelligence, including natural intelligence (human and animal), narrow AI, AGI, and ASI. We found that LLM model versions tend to be fragile and incremental as a result of memorisation only with progress likely driven by the size of training data. The results were compared with a hybrid neurosymbolic approach that theoretically guarantees universal intelligence based on the principles of algorithmic probability and Kolmogorov complexity. The method outperforms LLMs in a proof-of-concept on short binary sequences. We prove that compression is equivalent and directly proportional to a system’s predictive power and vice versa. That is, if a system can better predict it can better compress, and if it can better compress, then it can better predict. Our findings strengthen the suspicion regarding the fundamental limitations of LLMs, exposing them as systems optimised for the perception of mastery over human language.
Keywords: ARC tests, prediction, compression, program synthesis, inverse problems, causal AI, symbolic regression, comprehension, Superintelligence, Generative AI, symbolic computation, hybrid computation, Neurosymbolic computation.
@@ 1 Introduction
We are heavily biased to believe that the way humans think and act represents the acme of intelligence, even in instances where we may be limited, or flawed or irrational, or engaged in narrowly specific human (and often mundane) activities like chatting or washing dishes.
There will always be a natural tendency to overrate our own intelligence, to the detriment of efforts to devise a possibly more objective and quantitative measure of intelligence. But the question is exactly what that more objective test of intelligence might look like.
@@
One of the greatest realisations from the impressive apparent performance of Large Language Models (LLMs) is that language and other areas of human intellect may be overrated and are more dependent than we thought on memorisation and statistical pattern matching than critical thinking or other features of general intelligence.
One of the first metrics for intelligence was introduced by Charles Spearman in 1904 [1]. He proposed specific tests called ‘s’ that would each contribute to a general intelligence test under the name ‘g’, representing the common cognitive ability underlying performance in various mental tasks. Specific intelligences that contribute to the estimation of the g factor are verbal comprehension, perceptual reasoning, working memory, processing speed, quantitative reasoning, abstract reasoning, spatial ability, memory retrieval, auditory processing, and fluid reasoning. Some LLM benchmarks test for different factors, with several benchmarks based on correct answers versus hallucinations; some of which are very human-centric metrics related to human’s biological peculiarities and shared history.
A common psychological perspective sees intelligence through the lens of IQ tests, particularly the g-factor, a psychometric construct introduced by Spearman that quantifies the positive correlations between cognitive abilities. This framework is consistently linked to a human-centric perspective of what intelligence is and, therefore, biased towards circular reasoning. The concept of intelligence testing has been explored by researchers in different fields, including starting with machine intelligence rather than biological or human intelligence [2, 3, 4, 5, 6]. Some scholars argue that intelligence can be objectively defined through tests that evaluate specific computational abilities essential to demonstrate intelligent behaviour, rather than trying to define intelligence itself in absolute terms [2, 6, 3, 4]. This perspective shifts the focus from an abstract or philosophical definition to a practical, measurable framework assessing an entity’s capacity for problem-solving, pattern recognition, and adaptive learning within a structured system. It reflects an operational turn in the study of intelligence, emphasizing the design of formal benchmarks and quantifiable metrics. However, this approach is not without its philosophical challenges. By reducing intelligence to observable outputs, it risks overlooking the role of internal representation, consciousness, or semantic understanding–dimensions emphasized in critiques like Searle’s. In response to such concerns, researchers have sought to ground their metrics in more fundamental notions of computation and inference. For example, Gregory Chaitin [7] proposed that formal definitions of intelligence and its components should emerge from the mathematical theory of algorithmic complexity. Similarly, Solomonoff [8] advanced the idea of evaluating intelligence through algorithmic probability, laying the foundation for optimal prediction frameworks. These formal approaches, further developed in universal models like Hutter’s AIXI [9], attempt to reconcile objective evaluation with theoretical generality, but they still provoke debate regarding their ability to fully capture the qualitative essence of intelligence.
Based on these ideas, some tests for machine, human, and non-human entities have been proposed [10, 11, 4]. A generally accepted approach is that intelligence may be fundamentally linked to compression [5]—i.e., the ability to represent complex data in a simpler form while trying to lose the least information as possible. This suggests that intelligence involves identifying patterns, making predictions, and generating concise explanations for observed phenomena. Such an approach provides a unified framework for understanding both human and artificial intelligence, moving beyond traditional tests and philosophical debates to a measurable and practical foundation.

### v5

# Can Complexity and Uncomputability Explain Intelligence? SuperARC: A Test for Artificial Super Intelligence Based on Recursive Compression

@@ Abstract
We introduce an increasing-complexity, open-ended, and human-agnostic metric to evaluate foundational and frontier AI models in the context of Artificial General Intelligence (AGI) and Artificial Super Intelligence (ASI) claims. Unlike other tests that rely on human-centric questions and expected answers, or on pattern-matching methods, the test here introduced is grounded on fundamental mathematical areas of randomness and optimal inference. We argue that human-agnostic metrics based on the universal principles established by Algorithmic Information Theory (AIT) formally framing the concepts of model abstraction and prediction offer a powerful metrological framework. When applied to frontiers models, the leading LLMs outperform most others in multiple tasks, but they do not always do so with their latest model versions, which often regress and appear far from any global maximum or target estimated using the principles of AIT defining a Universal Intelligence (UAI) point and trend in the benchmarking. Conversely, a hybrid neuro-symbolic approach to UAI based on the same principles is shown to outperform frontier specialised prediction models in a simplified but relevant example related to compression-based model abstraction and sequence prediction. Finally, we prove and conclude that predictive power through arbitrary formal theories is directly proportional to compression over the algorithmic space, not the statistical space, and so further AI models’ progress can only be achieved in combination with symbolic approaches that LLMs developers are adopting often without acknowledgement or realisation.
Keywords: ARC-AGI test, prediction, recursive compression, program synthesis, inverse problems, causal AI, symbolic regression, comprehension, Superintelligence, Generative AI, symbolic computation, hybrid computation, Neurosymbolic computation.
@@ 1 Introduction
As we enter an AI test saturation phase, where all AI models claim to be the best or front runners on all available AI tests, new tests ideally orthogonal to current ones have to continue to be developed that can keep up with new models to keep increasingly challenging them.
Historically, humans have been heavily biased to believe that the way humans think and act represents the acme of intelligence, and there remains the philosophical and scientific question of the extent to which we can achieve more objective or less human-centric measures of intelligence.
@@
The impressive performance of Large Language Models (LLMs) as language processing and generation tools evinces that language and other areas of human intelligence may be overrated and can be, in fact, more dependent than we thought on aspects of memorisation and statistical pattern matching.
A common psychological perspective sees intelligence through the lens of IQ tests; one of the first was the g-factor, a psychometric construct introduced by Spearman [1] that quantifies the positive correlations between cognitive abilities. This framework is consistently linked to a human-centric perspective of what intelligence is and, therefore, biased towards circular reasoning. In the context of AI, some LLM benchmarks test for different factors, with several benchmarks based on correct answers versus hallucinations; some of which are also very human-centric metrics related to humans’ biological peculiarities and shared history.
Some scholars argue that intelligence can be objectively defined through tests that evaluate specific computational abilities essential to demonstrate intelligent behaviour, rather than trying to define intelligence itself in absolute terms [2, 3, 4, 5, 6]. This perspective shifts the focus from an abstract or philosophical definition to a practical, measurable framework assessing an entity’s capacity for problem-solving, pattern recognition, and adaptive learning within a structured system.
This reflects an operational turn in the study of intelligence, emphasising the design of formal benchmarks and quantifiable metrics. However, this approach is not without philosophical challenges. By reducing intelligence to observable outputs, it risks overlooking the role of internal representation, consciousness, or semantic understanding—dimensions emphasised in critiques like Searle’s.
An approach toward tackling such issues is to ground intelligence metrics in more fundamental notions of computation and mathematics. For example, the concepts of randomness, prediction and inference as defined by Gregory Chaitin, Andrey Kolmogorov, and Ray Solomonoff. Chaitin [7] proposed that formal definitions of intelligence and its components should be based on algorithmic complexity, a formal mathematical theory able to define the concept of randomness as opposed to intelligence. Similarly, Solomonoff [8] advanced the idea of evaluating intelligence through algorithmic probability, laying the foundation for optimal prediction frameworks (or universal “Bayesian” inference). These approaches further motivated approaches such as Hutter’s AIXI [9] in an attempt to reconcile objective evaluation with theoretical generality in the context of learning. Algorithmic complexity, algorithmic probability, and algorithmic randomness comprise the most important concepts in algorithmic information theory (AIT) [10, 11, 12, 13] and provide the accepted mathematical definitions of randomness and optimal inference (induction and abduction) going beyond simplistic statistical tests based on methods such as GZIP or LZW, popular pattern-matching compression methods that are more closely related to Shannon entropy than to model synthetization and predictive inference.
At recent public events, speaking about the foundations of AI and AGI, some leaders in the AI industry have drawn strong parallels between algorithmic complexity, data compression, and AI [14, 15]. Although these terminologies, such as AGI and ASI, are currently loosely defined in the scientific literature, these claims and the current understanding make the connection between LLMs (or any other generative AI), algorithmic complexity, and data compression clearer and more explicit, even calling it fundamental for general and super intelligence, artificial or natural.
Based on these arguments connecting intelligence to recursive compression [6], some tests for machine, human, and non-human entities have been proposed in [16, 17, 5]. Section 6.2 presents a reflection on that property of intelligence to involve the identification of recursive patterns, planning from prediction, and the generation of concise explanations for observed complex phenomena. Recursive compression here means the ability to represent an observation in a condensed manner by taking advantage of aspects of the data’s regularities beyond statistical pattern matching. This is, by selecting and keeping as many as possible the features that make the explanation executable and predictive of the explanandum future states.
Despite the interesting theoretical arguments that could be drawn from these connections, one argument is that they would only be valid under idealised conditions (unbounded data access/storage, perfect optimisation, appropriate inductive biases), which are rarely met in practice. As seen in real-world problems, even simple datasets with specific distributions can lead to optimisation toward local minima that do not correspond to minimal algorithmic descriptions.